7A 40kg mass is shot from a cannon at 50 m/s. The cannon recoils at 2 m/s. What is the mass of the cannon? d 105kg man is motionless on a frozen, frictionless pond. To get to the shore, he throws his 10ke a) A strande backpack at 2 m/s away from the shore. How fast does the man move? 8) b) If the pond is 30m away, how much time will it take for the man to get to the shore? A raft filled with the captain and gear has a mass of 3000kg is on the frozen, frictionless pond. The captain throws a 400kg box away from the shore at 0.5 m/s. 9) a) How fast does the raft move? b) How much time will the raft take to get to the shore 30m away? 10) A 5g bullet is shot into (and stays in) a 200g block of wood. If the wood+ bullet combo moves at 13 m/s, what is the velocity of the bullet as it strikes the wood? 11) A0.5kg superball is thrown at a student and knocks him in the head. If the 40kg student moves forward at 1.3 m/s and the superball bounces back at 20 m/s, then how fast was the ball thrown at the student?
Answer #1

10.

This is a case of completely inelastic collision:

Using momentum conservation in this case:

Pi = Pf

Mb*Ub + Mw*Uw = (Mb + Mw)*V

Ub = initial speed of bullet = ?

Uw = initial speed of wood = 0 m/sec

V = final speed of wood + bullet = 13 m/sec

Mb = 5 gm = 0.005 kg

Mw = 200 gm = 0.2 kg

So Using these values:

0.005*Ub + 0.2*0 = (0.005 + 0.2)*13

Ub = (0.005 + 0.2)*13/0.005 = 533 m/sec = Initial speed of bullet

11.

Again Using momentum conservation:

Pi = Pf

Mb*Ub + Ms*Us = Mb*Vb + Ms*Vs

Ub = initial speed of superball = ?

Us = initial speed of student = 0 m/sec

Vb = final speed of superball = -20 m/sec (See that ball bounces back, So direction will be opposite of initial direction)

Vs = final speed of student = +1.3 m/sec

Mb = 0.5 kg

Ms = 40 kg

Using these values:

0.5*Ub + 40*0 = 0.5*(-20) + 40*1.3

Ub = (0.5*(-20) + 40*1.3)/0.5

Ub = 84 m/sec = initial speed of superball

Please Ask rest of the questions as a new question, I will be happy to help. As HOMEWORKLIB RULES guideline I've answered one question. (In this case 2)

Please Upvote.